OpenStudy (anonymous):

[lnx]^7 FIND the derivative

OpenStudy (agent0smith):

Use the chain rule: \[\Large y = (\ln x) ^7\] Multiply by the power (7), reduce the power by one, then multiply by the derivative of the inside function (derivative of lnx is 1/x) \[\Large y \prime = 7 (\ln x) ^6 \left( \frac{ 1 }{ x } \right)\]

OpenStudy (anonymous):

do i have to simplify the ln(x) further

OpenStudy (agent0smith):

You can't. You can just move the 1/x \[\Large y \prime = \frac{ 7 }{ x } (\ln x) ^6 \]